# End-to-end Autonomous Driving: Challenges and Frontiers

### v1

@@ 4.1.2 Language as Input
Humans drive cars using both visual perception and intrinsic knowledge, such as traffic rules and desired routes, which together form causal behaviors. In some areas related to autonomous driving, such as robotics and indoor navigation (also called embodied AI), incorporating natural language as fine-grained instructions to control the visuomotor agent has achieved notable progress [147, 148, 149, 150].
However, the outdoor autonomous driving task owns different characteristics compared to indoor robotic applications in the following contexts: (1) The outdoor environment is unknown, and the vehicle cannot explore back and forth. (2) There are few distinctive anchor landmarks, posing great challenges in grounding language instructions. (3) Driving scenarios are much more complex, with a continuous action space and highly dynamic agents. Safety is a top priority during maneuvering. To incorporate linguistic knowledge into driving actions, the Talk2Car [151] dataset provides a benchmark for localizing referred objects in outdoor environments. The Talk2Nav [152], TouchDown [153] and Map2Seq [154] datasets introduce visual language navigation tasks using Google Street View. They model the world as a discrete connectivity graph and require navigating toward the goal in a node selection format. HAD [155] first takes human-to-vehicle advice and adds a visual grounding task with an LSTM-based controller. Sriram et al. [156] encode natural language instructions into high-level behaviors including turning left, turning right, not turning left, etc., and verify their language-guided navigation approach in the CARLA simulator. [157, 158] tackle the low-level real-time control problem by attending to the textual action command. Recently, CLIP-MC [159] and LM-Nav [160] utilize CLIP [161] which benefits from large-scale visual language pre-training to extract both linguistic knowledge from instructions and visual features from images. They showcase the advantages of the pre-trained model and provide an attractive prototype for solving complex navigation tasks with multi-modal models.
Despite the successful attempts to use CLIP for landmark feature extraction, the application of large language models (LLMs) such as GPT-3 [162] or instructive language models like ChatGPT [163, 164, 165] in the autonomous driving field remains unclear. Modern LLMs present more opportunities to handle sophisticated linguistic instructions. However, considering their long inference times and instability, it is also crucial to determine the interaction mode for on-road application. Moreover, current language-guided navigation works validate their effectiveness in simulation or specific robot embodiment, and a large-scale benchmark comprising meaningful linguistic prompts is missing.
4.2 Visual Abstraction
@@
An end-to-end autonomous driving system achieves the maneuvering task roughly in two stages: encoding the state space into a latent feature representation, and then decoding the driving policy with the intermediate features. In the case of urban driving, the input state, i.e., the surrounding environment and ego state, is much more diverse and high-dimensional compared to common policy learning benchmarks such as video games [18, 19, 166]. Therefore, it is helpful to first pre-train the visual encoder of the network using proxy pre-training tasks. This enables the network to effectively extract useful information for driving, thus facilitating the subsequent policy decoding stage while also meeting memory and model size constraints for all end-to-end algorithms. Furthermore, this can improve the sample efficiency for RL methods.
@@
The process of visual abstraction, or representation learning, often incorporates certain inductive biases or prior information. To achieve a more compact representation than raw images, some methods directly utilize semantic segmentation masks from pre-trained segmentation networks as the input representation for subsequent policy training [167, 168]. SESR [169] takes this a step further by encoding the segmentation masks into class-disentangled representations through a VAE [170]. In [171, 172], predicted affordance indicators, such as traffic light states, speed, offset to the lane center, hazard indicators, and distance to the leading vehicle, are used as the representation for policy learning.
@@
After observing that segmentation or affordances as a representation can create a bottleneck defined by humans and result in the loss of useful information, some have chosen intermediate latent features from pre-training tasks as effective representations. PIE-G [173] has demonstrated that early layers of the ImageNet [174] pre-trained model can serve as effective representations. [18, 19, 175, 176] employ latent representations pre-trained with tasks including semantic segmentation and/or affordance predictions as input for RL training and achieve excellent performance. In [177], latent features in VAE are augmented by attention maps obtained from the diffused boundary of segmentation and depth maps to highlight important regions. PPGeo [12] proposes to learn effective representation through motion prediction together with depth estimation in a self-supervised way on unlabeled driving videos. TARP [178] utilizes data from a series of previous tasks to perform different tasks-related prediction tasks to acquire useful representations. In [179], the latent representation is learned by approximating the $\pi$-bisimulation metric, which is comprised of differences of rewards and outputs from the dynamics model. Besides these pre-training tasks with supervised prediction, unsupervised contrastive learning based on augmented views is adopted in [133]. ACO [36] further adds steering angle discrimination into the contrastive learning structure.
As current methods mainly depend on human-defined pre-training tasks, there inevitably exist possible information bottlenecks in the learned representation, and redundant information unrelated to driving decisions may be included. Thus, how to better extract critical information for driving policy during the representation learning process remains an open question.
4.3 World Model and Model-based RL
@@
Deep reinforcement learning typically suffers from the challenge of high sample complexity, which is especially pronounced for tasks like autonomous driving due to the large size of the sample space. Model-based reinforcement learning (MBRL) offers a promising direction to improve sample efficiency by allowing agents to interact with the learned world model instead of the actual environment. MBRL methods explicitly model the world model/environment model, which is composed of transition dynamics and reward function, and the agent can interact with it at a low cost. This is particularly helpful in autonomous driving, as 3D simulators like CARLA are relatively slow.
@@
Modeling the highly complex and dynamic environment in driving is a challenging task. To simplify the problem, Chen et al. [20] assume that the world is on rails, factoring the transition dynamics into a non-reactive world model and a simple kinematic bicycle model for the ego vehicle. They enrich the labels of a static dataset by utilizing this factorized world model together with a reward function to optimize for better labels via dynamic programming. In [135], a probabilistic sequential latent model is used as the world model to reduce the sample complexity for RL learning. To address the potential inaccuracy of the learned world model, [180] uses an ensemble of multiple world models to provide uncertainty evaluation. Based on the uncertainty, the imaginary rollout between the world model and the policy agent could be truncated and adjusted accordingly. Motivated by the successful MBRL model Dreamer [83], ISO-Dream [181] considers non-deterministic factors in the environment and decouples visual dynamics into controllable and uncontrollable states. Then, the policy is trained on the disentangled states, considering the uncontrollable factors (such as the motion of other agents) explicitly.
@@
It is worth noting that learning the world model in raw image space is not suitable for autonomous driving. Important small details, such as traffic lights, would easily be missed in the predicted images. To tackle this, MILE [182] incorporates the world model in the BEV semantic segmentation space. It combines world modeling with imitation learning by adopting the Dreamer-style world model learning as an auxiliary task. SEM2 [134] also extends the Dreamer structure but with BEV segmentation maps, and uses RL for training. Besides directly using the learned world model for MBRL, DeRL [176] combines a model-free actor-critic framework with the world model. Specifically, the learned world model provides self-assessments of the current action, which is combined with the state value from the critic to better access the actor’s performance.
@@
World model learning or MBRL for end-to-end autonomous driving is an emerging and promising direction as it greatly reduces the sample complexity for RL, and understanding the world is helpful for driving. However, as the driving environment is highly complex and dynamic, further study is still needed to determine what needs to be modeled and how to model the world effectively.
4.4 Multi-task Learning with Policy Prediction
@@
Multi-task learning (MTL) involves jointly performing several related tasks based on a shared representation through separate branches/heads. MTL offers significant computational cost reduction by using a single model for multiple tasks. Additionally, related domain knowledge is shared within the shared model, and the task relationship can be better exploited to improve the generalization ability and robustness of the model [183, 184]. Consequently, MTL is a good fit for end-to-end autonomous driving, where the ultimate policy prediction requires a comprehensive understanding of the current environment.
@@
In contrast to common vision tasks where dense predictions are needed, end-to-end autonomous driving predicts a sparse signal. The sparse supervision here brings a challenge for the input encoder to extract useful information for decision-making. For image input, auxiliary tasks such as semantic segmentation [31, 28, 136, 185, 186, 187] and depth estimation [31, 28, 185, 186, 187] are commonly adopted in end-to-end autonomous driving models. Semantic segmentation ensures that the model gains a high-level understanding of the scene and recognizes objects of different categories; depth estimation enables the model to understand the 3D geometry of the environment and better estimate distances to critical objects. By performing these tasks, the image encoder can better extract useful and meaningful feature representations for subsequent planning. Besides auxiliary tasks on perspective images, 3D object detection [31, 28, 52] is also useful for the LiDAR encoder. As BEV becomes a natural and popular representation for autonomous driving, tasks such as HD map mapping and BEV segmentation are included in models [31, 28, 11, 52, 29, 30, 23, 144] that aggregate features in BEV space. Moreover, in addition to these vision tasks as multi-tasks, [29, 188, 185] also predict visual affordances including traffic light states, distance to the junction, and distances to opposite lanes, etc.
Multi-task learning for end-to-end autonomous driving has demonstrated its effectiveness in improving performance and providing interpretability of the autonomous driving model. However, the optimal combination of auxiliary tasks and the appropriate weighting of their losses to achieve the best performance remain to be explored. Additionally, constructing large-scale datasets with multiple types of aligned and high-quality annotations presents a significant challenge.
@@
4.5 Policy Distillation
@@
As imitation learning, or its predominant sub-category, behavior cloning, is simply supervised learning that mimics expert behaviors, corresponding methods usually follow the “Teacher-Student” paradigm. Teachers, such as the handcrafted expert autopilot provided by CARLA, have access to the ground-truth states of surrounding agents and map elements, while students are directly supervised by the collected expert trajectory or control signals with raw sensor input only. This raises great challenges for student models, as they must not only extract perceptual features but also learn a driving policy from scratch.
@@
To address the difficulties above, a few studies propose to divide the learning process into two stages, i.e., training a teacher network and then distilling the policy to the ultimate student network. In particular, Chen et al. [5] first employ a privileged agent to learn how to act with direct access to the state of the environment. They then let the sensorimotor agent (student network) closely imitate the privileged agent, with distillation occurring at the output stage. With more compact BEV representations as input for the privileged agent, it provides stronger generalization abilities and supervision than the original expert. The process is depicted in Fig. 5. LAV [52] further empowers the privileged agent to predict the trajectories of all nearby vehicles and distills this ability to the student network which uses vision features.
@@
Apart from directly supervising planning results, several works train their predictive models by distilling knowledge at the feature level. For example, FM-Net [189] employs off-the-shelf networks, including segmentation and optical flow models, as auxiliary teachers to guide feature training. SAM [190] adds L2 feature loss between the teacher and student networks, where the teacher network predicts control signals from ground-truth semantic segmentation maps, and stop intention values. WoR [20] learns a model-based action-value function and then uses it to supervise the visuomotor policy. CaT [23] recently introduces BEV safety hints in the IL-based privileged expert training and distills in BEV space to align features. Roach [21] proposes to train a stronger privileged expert with RL, eliminating the upper bound of imitation learning. It incorporates multiple distillation targets, i.e., action distribution prediction, value estimation, and latent features. By leveraging the powerful RL expert, TCP [22] achieves a new state-of-the-art on the CARLA leaderboard with a single camera as visual input.
@@
Though a great amount of effort has been devoted to designing a more robust expert and transmitting knowledge from teachers to students at various levels, the teacher-student paradigm still suffers from inefficient distillation. As shown in all previous works, the visuomotor networks exhibit large performance gaps compared to their privileged agents. For instance, the privileged agent has access to the ground-truth states of traffic lights, which are small objects in images and pose a challenge to distill corresponding features. This may lead to causal confusion for students, as discussed in Sec. 4.7. Thus, it is worth exploring how to draw more inspiration from general distillation methods in machine learning to minimize the gap.
@@
Figure 6: Summary of the different forms of interpretability. They aid in human comprehension of the decision-making processes of end-to-end models, perception failures, and the reliability of the outputs.
4.6 Interpretability
@@
Interpretability plays a critical role in autonomous driving [191]. It enables engineers and researchers to better test, debug, and improve the system, provides performance guarantees from a societal perspective, increases users’ trust, and promotes public acceptance. However, achieving interpretability in end-to-end autonomous driving models, which are often referred to as black boxes, is challenging.
@@
Given a trained autonomous driving model, some post-hoc X-AI (explainable AI) techniques could be applied to the learned model to gain saliency maps [192, 193, 194, 195, 185]. Saliency maps highlight specific regions in the visual input on which the model primarily relies for planning. However, this approach provides limited information, and its effectiveness and validity are difficult to evaluate. Instead, we focus on autonomous driving frameworks that directly enhance interpretability in their model design. We will introduce each category of interpretability in Fig. 6 below.
@@
Attention Visualization: The attention mechanism usually provides a certain degree of interpretability. In [195, 196, 185, 188, 33], a learned attention weight is applied to aggregate important features from the intermediate feature map. Attention weights are learned to adaptively combine ROI pooled features from different object regions [197] or a fixed grid [198]. NEAT [11] iteratively aggregates features to predict attention weights and refine the aggregated feature. Recently, the Transformer attention mechanism has become commonly used in many autonomous driving models. Transformer attention blocks are employed in [199, 6, 28, 31, 29, 143, 11] to better aggregate the information from sensor inputs, and attention maps display important regions in the input for driving decisions. In PlanT [200], the attention layers process features from different vehicles, providing interpretable insights into the corresponding action. Similar to post-hoc saliency methods, although the learned attention maps could offer some straightforward clues about the models’ focus, their faithfulness and utility remain limited.
@@
Interpretable Tasks: In a deep driving model, the inputs are initially encoded into intermediate representations for subsequent predictions. Therefore, many IL-based works introduce interpretability by decoding the latent feature representations into other meaningful information, such as semantic segmentation [2, 29, 52, 11, 28, 31, 159, 185, 15, 136, 186, 187], depth estimation [28, 31, 185, 15, 186], object detection [2, 52, 28, 31], affordance predictions [29, 185, 188], motion prediction [2, 52], and gaze map estimation [201]. Although these methods provide interpretable information, most of them only treat these predictions as auxiliary tasks [28, 31, 11, 185, 15, 136, 186, 188], with no explicit impact on the final driving decision. Some [29, 52] do use these outputs for final driving actions, but they are incorporated solely for performing an additional safety check.
@@
Cost Learning: As discussed in Sec. 2.1.2, cost learning-based methods share some similarities with traditional modular autonomous driving systems and thus exhibit a certain level of interpretability. NMP [32] and DSDNet [202] construct the cost volume in conjunction with detection and motion prediction results. P3 [39] combines predicted semantic occupancy maps with comfort and traffic rules constraints to construct the cost function. Various representations, such as probabilistic occupancy and temporal motion fields [1], emergent occupancy [71], and freespace [70], are employed to score sampled trajectories. Factors including safety, comfort, traffic rules, and routes based on the perception and prediction outputs are explicitly included [203, 38] to construct the cost volumes.
@@
Linguistic Explainability: As one aspect of interpretability is to help humans understand the system, natural language is a suitable choice for this purpose. Kim et al. [33] generate the BDD-X dataset, pairing driving videos with descriptions and explanations. They also propose an autonomous driving model with a vehicle controller and an explanation generator, and force spatial attention weights of the two modules to be aligned. BEEF [204] proposes an explanation module that fuses the predicted trajectory and the intermediate perception features to predict justifications for the decision. In [205], a dataset named BBD-OIA is introduced, which includes annotations for both driving decisions and explanations for high-density traffic scenes. Recently, ADAPT [206] proposes a Transformer-based network to jointly estimate action, narration, and reasoning based on driving videos from the BBD-X dataset. Given the recent progress of multi-modality and foundation models, we believe that further combining language with autonomous driving models holds promise for achieving superior interpretability and performance, as discussed in Sec. 4.1.2.
@@
Uncertainty Modeling: Uncertainty is a quantitative approach for interpreting the dependability of model outputs. Since planning results are not always accurate or optimal, it is essential for designers and users to identify uncertain cases for improvement or necessary intervention. For deep learning, there are two types of uncertainty: aleatoric uncertainty and epistemic uncertainty. Aleatoric uncertainty is inherent to the task, while epistemic uncertainty is due to limited data or modeling capacity. In [207], quantitative evaluations of uncertainty for end-to-end autonomous driving systems are conducted, leveraging certain stochastic regularizations in the model to perform multiple forward passes as samples to measure the uncertainty. However, the requirement of multiple forward passes is not feasible in real-time scenarios. RIP [208] proposes capturing epistemic uncertainty with an ensemble of expert likelihood models and aggregating the results to perform safe planning. Regarding methods modeling aleatoric uncertainty, driving actions/planning and uncertainty (usually represented by variance) are explicitly predicted in [209, 210, 142]. With the predicted uncertainty, [209] chooses the output with the lowest uncertainty from multiple outputs, while [142] generates a weighted combination of proposed actions. VTGNet [210] does not directly use the uncertainty for planning but demonstrates that modeling data uncertainty can improve general performance. Currently, predicted uncertainty is mainly utilized in combination with hard-coded rules. Exploring better ways to model and utilize uncertainty for autonomous driving is necessary.

### v2

@@ 4.1.2 Language as Input
Humans drive using both visual perception and intrinsic knowledge which together form causal behaviors. In areas related to autonomous driving such as embodied AI, incorporating natural language as fine-grained knowledge and instructions to control the visuomotor agent has achieved notable progress [151, 152, 153, 154]. However, compared to robotic applications, the driving task is more straightforward without the need for task decomposition, and the outdoor environment is much more complex with highly dynamic agents but few distinctive anchors for grounding.
To incorporate linguistic knowledge into driving, a few datasets are proposed to benchmark outdoor grounding and visual language navigation tasks [155, 156, 157, 158]. HAD [159] takes human-to-vehicle advice and adds a visual grounding task. Sriram et al. [160] translate natural language instructions into high-level behaviors, while [161, 162] directly ground the texts. CLIP-MC [163] and LM-Nav [164] utilize CLIP [165] to extract both linguistic knowledge from instructions and visual features from images.
Recently, observing the rapid development of large language models (LLMs) [166, 167], works encode the perceived scene into tokens and prompt them to LLMs for control prediction and text-based explanations [168, 169, 170]. Researchers also formulate the driving task as a question-answering problem and construct corresponding benchmarks [171, 172]. They highlight that LLMs offer opportunities to handle sophisticated instructions and generalize to different data domains, which share similar advantages to applications in robotic areas [173]. However, LLMs for on-road driving could be challenging at present, considering their long inference time, low quantitative accuracy, and instability of outputs. Potential resolutions could be employing LLMs on the cloud specifically for complex scenarios and using them solely for high-level behavior prediction.
4.2 Dependence on Visual Abstraction
@@
End-to-end autonomous driving systems roughly have two stages: encoding the state into a latent feature representation, and then decoding the driving policy with intermediate features. In urban driving, the input state, i.e., the surrounding environment and ego state, is much more diverse and high-dimensional compared to common policy learning benchmarks such as video games [18, 174], which might lead to the misalignment between representations and necessary attention areas for policy making. Hence, it is helpful to design “good” intermediate perception representations, or first pre-train visual encoders using proxy tasks. This enables the network to extract useful information for driving effectively, thus facilitating the subsequent policy stage. Furthermore, this can improve the sample efficiency for RL methods.
4.2.1 Representation Design
Naive representations are extracted with various backbones. Classic convolutional neural networks (CNNs) still dominate, with advantages in translation equivariance and high efficiency [175]. Depth-pre-trained CNNs [176] significantly boost perception and downstream performance. In contrast, Transformer-based feature extractors [177, 178] show great scalability in perception tasks while not being widely adopted for end-to-end driving yet. For driving-specific representations, researchers introduce the concept of bird’s-eye-view (BEV), fusing different sensor modalities and temporal information within a unified 3D space [179, 180, 132, 131]. It also facilitates easy adaptions to downstream tasks [2, 30, 181]. In addition, grid-based 3D occupancy is developed to capture irregular objects and used for collision avoidance in planning [182]. Nevertheless, the dense representation brings huge computation costs compared to BEV methods.
Another unsettled problem is representations of the map. Traditional autonomous driving relies on HD Maps. Due to the high cost of availability of HD Maps, online mapping methods have been devised with different formulations, such as BEV segmentation [183], vectorized lanlines [184], centerlines and their topology [185, 186], and lane segments [187]. However, the most suitable formulation for end-to-end systems remains unvalidated.
Though various representation designs offer possibilities of how to design the subsequent decision-making process, they also place challenges as co-designing both parts is necessary for a whole framework. Besides, given the trends observed in several simple yet effective approaches with scaling up training resources [22, 28], the ultimate necessity of explicit representations such as maps is uncertain.
4.2.2 Representation Learning
Representation learning often incorporates certain inductive biases or prior information. There inevitably exist possible information bottlenecks in the learned representation, and redundant context unrelated to decisions may be removed.
@@ 4.2.2 Representation Learning
Some early methods directly utilize semantic segmentation masks from off-the-shelf networks as the input representation for subsequent policy training [188, 189]. SESR [190] further encodes segmentation masks into class-disentangled representations through a VAE [191]. In [192, 193], predicted affordance indicators, such as traffic light states, offset to the lane center, and distance to the leading vehicle, are used as representations for policy learning.
@@
Observing that results like segmentation as representations can create bottlenecks defined by humans and result in loss of useful information, some have chosen intermediate features from pre-training tasks as effective representations for RL training [18, 19, 194, 195]. In [196], latent features in VAE are augmented by attention maps obtained from the diffused boundary of segmentation and depth maps to highlight important regions. TARP [197] utilizes data from a series of previous tasks to perform different tasks-related prediction tasks to acquire useful representations. In [198], the latent representation is learned by approximating the $\pi$-bisimulation metric, which is comprised of differences of rewards and outputs from the dynamics model. ACO [36] learns discriminative features by adding steering angle categorization into the contrastive learning structure. Recently, PPGeo [12] proposes to learn effective representation through motion prediction together with depth estimation in a self-supervised way on uncalibrated driving videos. ViDAR [199] utilizes the raw image-point cloud pairs and pretrains the visual encoder with a point cloud forecasting pre-task. These works demonstrate that self-supervised representation learning from large-scale unlabeled data for policy learning is promising and worthy of future exploration.
4.3 Complexity of World Modeling for Model-based RL
@@
Deep RL typically suffers from the high sample complexity, which is pronounced in autonomous driving. Model-based reinforcement learning (MBRL) offers a promising direction to improve sample efficiency by allowing agents to interact with the learned world model instead of the actual environment. MBRL methods employ an explicit world (environment) model, which is composed of transition dynamics and reward functions. This is particularly helpful in driving, as simulators like CARLA are relatively slow.
@@
However, modeling the highly dynamic environment is a challenging task. To simplify the problem, Chen et al. [20] factor the transition dynamics into a non-reactive world model and a simple kinematic bicycle model. In [137], a probabilistic sequential latent model is used as the world model. To address the potential inaccuracy of the learned world model, Henaff et al. [200] train the policy network with dropout regularization to estimate the uncertainty cost. Another approach [201] uses an ensemble of multiple world models to provide uncertainty estimation, based on which imaginary rollouts could be truncated and adjusted accordingly. Motivated by Dreamer [82], ISO-Dream [202] decouples visual dynamics into controllable and uncontrollable states, and trains the policy on the disentangled states.
@@
It is worth noting that learning world models in raw image space is non-trivial for autonomous driving. Important small details, such as traffic lights, would easily be missed in predicted images. To tackle this, GenAD [203] and DriveWM [204] employ the prevailing diffusion technique [205]. MILE [206] incorporates the Dreamer-style world model learning in the BEV segmentation space as an auxiliary task besides imitation learning. SEM2 [136] also extends the Dreamer structure but with BEV map inputs, and uses RL for training. Besides directly using the learned world model for MBRL, DeRL [195] combines a model-free actor-critic framework with the world model, by fusing self-assessments of the action or state from both models.
@@
World model learning for end-to-end autonomous driving is an emerging and promising direction as it greatly reduces the sample complexity for RL, and understanding the world is helpful for driving. However, as the driving environment is highly complex and dynamic, further study is still needed to determine what needs to be modeled and how to model the world effectively.
4.4 Reliance on Multi-Task Learning
@@
Multi-task learning (MTL) involves jointly performing several related tasks based on a shared representation through separate heads. MTL provides advantages such as computational cost reduction, the sharing of relevant domain knowledge, and the ability to exploit task relationships to improve model’s generalization ability [207]. Consequently, MTL is well-suited for end-to-end driving, where the ultimate policy prediction requires a comprehensive understanding of the environment. However, the optimal combination of auxiliary tasks and appropriate weighting of losses to achieve the best performance presents a significant challenge.
@@
In contrast to common vision tasks where dense predictions are closely correlated, end-to-end driving predicts a sparse signal. The sparse supervision increases the difficulty of extracting useful information for decision-making in the encoder. For image input, auxiliary tasks such as semantic segmentation [31, 28, 139, 208, 209, 210] and depth estimation [31, 28, 208, 209, 210] are commonly adopted in end-to-end autonomous driving models. Semantic segmentation helps the model gain a high-level understanding of the scene; depth estimation enables the model to capture the 3D geometry of the environment and better estimate distances to critical objects. Besides auxiliary tasks on perspective images, 3D object detection [31, 28, 52] is also useful for LiDAR encoders. As BEV becomes a natural and popular representation for autonomous driving, tasks such as BEV segmentation are included in models [31, 28, 11, 52, 29, 30, 23, 148] that aggregate features in BEV space. Moreover, in addition to these vision tasks, [29, 211, 208] also predict visual affordances including traffic light states, distances to opposite lanes, etc. Nonetheless, constructing large-scale datasets with multiple types of aligned and high-quality annotations is non-trivaial for real-world applications, which remain as a great concern due to current models’ reliance on MTL.
@@
4.5 Inefficient Experts and Policy Distillation
@@
As imitation learning, or its predominant sub-category, behavior cloning, is simply supervised learning that mimics expert behaviors, corresponding methods usually follow the “Teacher-Student” paradigm. There lie two main challenges: (1) Teachers, such as the handcrafted expert autopilot provided by CARLA, are not perfect drivers, though having access to ground-truth states of surrounding agents and maps. (2) Students are supervised by the recorded output with sensor input only, requiring them to extract perceptual features and learn policy from scratch simultaneously.
@@
A few studies propose to divide the learning process into two stages, i.e., training a stronger teacher network and then distilling the policy to the student. In particular, Chen et al. [5, 52] first employ a privileged agent to learn how to act with access to the state of the environment, then let the sensorimotor agent (student) closely imitate the privileged agent with distillation at the output stage. More compact BEV representations as input for the privileged agent provide stronger generalization abilities and supervision than the original expert. The process is depicted in Fig. 5.
@@
Apart from solely supervising planning results, several works also distill knowledge at the feature level. For example, FM-Net [212] employs segmentation and optical flow models as auxiliary teachers to guide feature training. SAM [213] adds L2 feature loss between teacher and student networks, while CaT [23] aligns features in BEV. WoR [20] learns a model-based action-value function and then uses it to supervise the visuomotor policy. Roach [21] trains a stronger privileged expert with RL, eliminating the upper bound of BC. It incorporates multiple distillation targets, i.e., action distribution, values/rewards, and latent features. By leveraging the powerful RL expert, TCP [22] achieves a new state-of-the-art on the CARLA leaderboard with a single camera as visual input. DriveAdpater [181] learns a perception-only student and adapters with the feature alignment objective. The decoupled paradigm fully enjoys the teacher’s knowledge and student’s training efficiency.
@@
Though huge efforts have been devoted to designing a robust expert and transferring knowledge at various levels, the teacher-student paradigm still suffers from inefficient distillation. For instance, the privileged agent has access to ground-truth states of traffic lights, which are small objects in images and thus hard to distill corresponding features. As a result, the visuomotor agents exhibit large performance gaps compared to their privileged agents. It may also lead to causal confusion for students (see Sec. 4.8). It is worth exploring how to draw more inspiration from general distillation methods in machine learning to minimize the gap.
@@
Figure 6: Summary of the different forms of interpretability. They aid in human comprehension of decision-making processes of end-to-end models and the reliability of outputs.
4.6 Lack of Interpretability
@@
Interpretability plays a critical role in autonomous driving [214]. It enables engineers to better debug the system, provides performance guarantees from a societal perspective, and promotes public acceptance. Achieving interpretability for end-to-end driving models, which are often referred to as “black boxes”, is more essential and challenging.
@@
Given trained models, some post-hoc X-AI (explainable AI) techniques could be applied to gain saliency maps [215, 216, 217, 218, 208]. Saliency maps highlight specific regions in the visual input on which the model primarily relies for planning. However, this approach provides limited information, and its effectiveness and validity are difficult to evaluate. Instead, we focus on end-to-end frameworks that directly enhance interpretability in their model design. We introduce each category of interpretability in Fig. 6 below.
@@
Attention Visualization: The attention mechanism provides a certain degree of interpretability. In [218, 219, 208, 211, 33], a learned attention weight is applied to aggregate important features from intermediate feature maps. Attention weights can also adaptively combine ROI pooled features from different object regions [220] or a fixed grid [221]. NEAT [11] iteratively aggregates features to predict attention weights and refine the aggregated feature. Recently, Transformer attention blocks are employed to better fuse different sensor inputs, and attention maps display important regions in the input for driving decisions [222, 28, 31, 29, 147]. In PlanT [223], attention layers process features from different vehicles, providing interpretable insights into the corresponding action. Similar to post-hoc saliency methods, although attention maps offer straightforward clues about models’ focus, their faithfulness and utility remain limited.
@@
Interpretable Tasks: Many IL-based works introduce interpretability by decoding the latent feature representations into other meaningful information besides policy prediction, such as semantic segmentation [2, 29, 52, 11, 28, 31, 163, 208, 15, 139, 209, 210, 224], depth estimation [28, 31, 208, 15, 209], object detection [2, 52, 28, 31], affordance predictions [29, 208, 211], motion prediction [2, 52], and gaze map estimation [225]. Although these methods provide interpretable information, most of them only treat these predictions as auxiliary tasks [28, 31, 11, 208, 15, 139, 209, 211], with no explicit impact on final driving decisions. Some [29, 52] do use these outputs for final actions, but they are incorporated solely for performing an additional safety check.
@@
Rules Integration and Cost Learning: As discussed in Sec. 2.1.2, cost learning-based methods share similarities with traditional modular systems and thus exhibit a certain level of interpretability. NMP [32] and DSDNet [226] construct the cost volume in conjunction with detection and motion prediction results. P3 [39] combines predicted semantic occupancy maps with comfort and traffic rules constraints to construct the cost function. Various representations, such as probabilistic occupancy and temporal motion fields [1], emergent occupancy [71], and freespace [70], are employed to score sampled trajectories. In [227, 38, 125], human expertise and pre-defined rules including safety, comfort, traffic rules, and routes based on perception and prediction outputs are explicitly included to form the cost for trajectory scoring, demonstrating improved robustness and safety.
@@
Linguistic Explainability: As one aspect of interpretability is to help humans understand the system, natural language is a suitable choice for this purpose. Kim et al. [33] and Xu et al. [228] develop datasets pairing driving videos or images with descriptions and explanations, and propose end-to-end models with both control and explanation outputs. BEEF [229] fuses the predicted trajectory and the intermediate perception features to predict justifications for the decision. ADAPT [230] proposes a Transformer-based network to jointly estimate action, narration, and reasoning. Recently, [171, 172, 169] resort to the progress of multi-modality and foundation models, using LLMs/VLMs to provide decision-related explanations, as discussed in Sec. 4.1.2.
@@
Uncertainty Modeling: Uncertainty is a quantitative approach for interpreting the dependability of deep learning model outputs [231, 232], which can be helpful for designers and users to identify uncertain cases for improvement or necessary intervention. For deep learning, there are two types of uncertainty: aleatoric uncertainty and epistemic uncertainty. Aleatoric uncertainty is inherent to the task, while epistemic uncertainty is due to limited data or modeling capacity. In [233], authors leverage certain stochastic regularizations in the model to perform multiple forward passes as samples to measure the uncertainty. However, the requirement of multiple forward passes is not feasible in real-time scenarios. Loquercio et al. [232] and Filos et al. [234] propose capturing epistemic uncertainty with an ensemble of expert likelihood models and aggregating the results to perform safe planning. Regarding methods modeling aleatoric uncertainty, driving actions/planning and uncertainty (usually represented by variance) are explicitly predicted in [235, 236, 146]. Such methods directly model and quantify the uncertainty at the action level as a variable for the network to predict. The planner would generate the final action based on the predicted uncertainty, either choosing the action with the lowest uncertainty from multiple actions [235] or generating a weighted combination of proposed actions based on the uncertainties [146]. Currently, predicted uncertainty is mainly utilized in combination with hard-coded rules. Exploring better ways to model and utilize uncertainty for autonomous driving is necessary.
4.7 Lack of Safety Guarantees
Ensuring safety is of utmost importance when deploying autonomous driving systems in real-world scenarios. However, the learning-based nature of end-to-end frameworks inherently lacks precise mathematical guarantees regarding safety, unlike traditional rule-based approaches [237].
Nevertheless, it should be noted that modular driving stacks have already incorporated specific safety-related constraints or optimizations within their motion planning or speed prediction modules to enforce safety [238, 239, 240]. These mechanisms can potentially be adapted for integration into end-to-end models as post-process steps or safety checks, thereby providing additional safety guarantees. Furthermore, the intermediate interpretability predictions, as discussed in Sec. 4.6, such as detection and motion prediction results, can be utilized in post-processing procedures.